OpenStudy (anonymous):

Hi I am helping my little sister with her math homework so if you could help me much would be appreciated. Thank you Compare and Contrast: Below are two equations. Solve each equation and compare the two solutions. Choose the statement that is true about each solution. Equation #1 Equation #2 6 - x = 5x + 30 4x - 8 = 3x + 2

OpenStudy (ash2326):

hey kari post the statements also

OpenStudy (anonymous):

None of the statements above describes solutions to equations shown. The solution to equation #1 is smaller than the solution to equation #2. The solution to equation #1 is the same as the solution to equation #2. The solution to equation #1 is larger than the solution to equation #2.

OpenStudy (ash2326):

solving the first equation 6-x=5x+30 bringing the x terms together -6x=30-6 -6x=24 x=-6 now solving the second equation 4x-8=3x+2 x=10 so solution of first one is smaller than the second one's solution second statement is the correct choice

OpenStudy (ash2326):

sorry i made a mistake x=-4 in first one , answer remains unchanged , i apologise

OpenStudy (anonymous):

Its all good so its the second statement. Thanks. Your really nice. I will give you a medal. (:

OpenStudy (ash2326):

thanks kary:)

OpenStudy (anonymous):

Its Kairi but no problem Ash

OpenStudy (ash2326):

sorry kairi:)